# Comparing diversity, negativity, and stereotypes in Chinese-language AI technologies: a case study on Baidu, Ernie and Qwen

### v1

@@ Abstract
Large Language Models (LLMs) and search engines have the potential to perpetuate biases and stereotypes by amplifying existing prejudices in their training data and algorithmic processes, thereby influencing public perception and decision-making. While most work has focused on Western-centric AI technologies, we study Chinese-based tools by investigating social biases embedded in the major Chinese search engine, Baidu, and two leading LLMs, Ernie and Qwen. Leveraging a dataset of 240 social groups across 13 categories describing Chinese society, we collect over 30k views encoded in the aforementioned tools by prompting them for candidate words describing such groups. We find that language models exhibit a larger variety of embedded views compared to the search engine, although Baidu and Qwen generate negative content more often than Ernie. We also find a moderate prevalence of stereotypes embedded in the language models, many of which potentially promote offensive and derogatory views. Our work highlights the importance of promoting fairness and inclusivity in AI technologies with a global perspective.
@@ 1 Introduction
The advent of novel Large Language Models (LLMs) has revolutionized the field of natural language processing (NLP), offering unprecedented potential for a wide range of applications, from machine translation and sentiment analysis to conversational agents and content generation (Zhu et al.,, 2024; Zhang et al.,, 2024; Deng et al.,, 2023). LLMs, such as GPT-4 and LLaMa, have demonstrated an impressive ability to understand and generate human-like text, making them invaluable tools in both academic research and commercial settings (Ziems et al.,, 2024; Minaee et al.,, 2024). However, along with their potential, these models also carry significant risks. As they become increasingly integrated into various applications, concerns about their reliability, ethical use, and potential for misuse have grown (May et al.,, 2019; Nadeem et al.,, 2021; Choenni et al.,, 2021; Navigli et al.,, 2023; Busker et al.,, 2023; Deshpande et al.,, 2023). In addition, the opaque nature of these models often makes it difficult to understand their inference processes, raising questions about accountability and transparency (Bender et al.,, 2021).
@@
One pressing issue associated with this technology is the presence of biases and stereotypes embedded within language models, which can stem from the data they are trained on, thus reflecting the prejudices and inequalities prevalent in broader cultural and social contexts (Liang et al.,, 2020). These biases can manifest in harmful ways, such as reinforcing stereotypes or supporting discriminatory decisions, thus perpetuating social injustices (Weidinger et al.,, 2021). Researchers have developed various methods to measure and mitigate these biases, proposing datasets and metrics to evaluate the models’ fairness (Zmigrod et al.,, 2019; Webster et al.,, 2020; Liang et al.,, 2020). Despite these efforts, significant challenges remain, particularly when addressing proprietary models where access to internal mechanisms remains restricted (Huang et al.,, 2024). This limitation hampers the effectiveness of many bias detection and mitigation techniques, necessitating further innovation and scrutiny in this area.
@@
Most research on bias and fairness in LLMs has focused on English and other widely spoken languages, leaving a gap in our understanding of how these models perform in different linguistic and cultural contexts (Ducel et al.,, 2023; Yang et al.,, 2024). Chinese, natively spoken by over 1.35 billion individuals ($\sim 17\%$ of the global population), has received comparatively little attention from the scientific community (Xu et al.,, 2024). This oversight is critical, given the unique cultural, social, and linguistic characteristics of the Chinese language. Addressing this gap is essential for developing fair and unbiased LLMs that can serve diverse populations effectively and ethically (Zhao et al.,, 2023; Huang and Xiong,, 2024). Consequently, our study aims to investigate the biases present in two leading Chinese LLMs, namely Ernie and Qwen, as well as in Baidu, the most popular search engine used in China; we include the latter following previous work (Choenni et al.,, 2021; Busker et al.,, 2023; Liu et al.,, 2024) that highlighted the relevance of online search engines at potentially perpetuating cultural biases and stereotypes, as well as pointing out how LLMs are being employed by online users to answer a variety of queries. Leveraging a dataset of social groups and categories pertaining to Chinese society, we use an empirical approach to study the social views encoded in the models. Following (Busker et al.,, 2023), we probe the models with varied questions, with the goal of eliciting the views that are encoded about different social groups. We then study the completions obtained in terms of diversity, negativity, and alignment. We use this approach as the internal models’ probability for predicted tokens is not available, and one cannot apply existing metrics for measuring biases associated with certain words (Kurita et al.,, 2019). Our work contributes to the literature on fairness and inclusivity in AI-generated content, with a discussion of the implications of our findings in the concluding section of the paper.
@@
The outline of the manuscript is the following: in the next section we overview related work; then, we describe the data collection and methodologies employed in our analyses; next, we present the results of our analyses; finally, we discuss our findings and their implications, mention limitations to our work and draw conclusions.
@@ 2 Related Work
In recent years, numerous studies Caliskan et al., (2017); Zhao et al., (2018); May et al., (2019); Kurita et al., (2019); Liang et al., (2020); Barikeri et al., (2021); Wald and Pfahler, (2023) have introduced various datasets and metrics to quantify social bias and stereotypes in pre-trained LLMs. However, these efforts face challenges with most proprietary models, as access to internal model information (e.g., probability of predicted tokens and internal embeddings) is restricted (Huang et al.,, 2023), thereby limiting the application of metrics such as the Log Probability Bias Score (LPBS) (Kurita et al.,, 2019) and the Sentence Embedding Association Test (SEAT) (May et al.,, 2019). In the following, we focus on contributions more directly related to our work and refer readers to  Gallegos et al., (2024) for a more comprehensive review of the literature on this topic.
@@
Choenni et al., (2021) introduced a method for eliciting stereotypes by retrieving salient attributes about social groups through the auto-completion functionality in search engines such as Google, Yahoo, and DuckDuckGo. They then analyzed the predictions of pre-trained language models for these attributes to assess how many stereotypes were encoded within the models. Their study also highlighted the evolution of stereotypes with modifications in training data during model fine-tuning stages. Building on this work, Busker et al., (2023) conducted an empirical study to explore stereotypical behavior in ChatGPT, focusing on U.S.-centric social groups. They presented stereotypical prompts in six formats (e.g., questions and statements) and across nine different social group categories (e.g., age, country, and profession). For each prompt, they asked ChatGPT to fill in a masked word and mapped the completions to sentiment levels to measure stereotypical behavior.
@@
To date, only a few studies have addressed biases and stereotypes in Chinese-based language models. Huang and Xiong, (2024) extended the Biases Benchmark (BBQ) dataset Parrish et al., (2022) from U.S. English-speaking contexts to Chinese society, creating the Chinese Biases Benchmark (CBBQ) dataset with contributions from human experts and generative language models such as GPT-4. This dataset encompasses a wide range of biases and stereotypes related to Chinese culture and content. Similarly, Zhao et al., (2023) introduced the CHBias dataset, which includes biases related to ageism, appearance, and other factors, for evaluating and mitigating biases in Chinese conversational models like CDial-GPT Wang et al., (2020) and EVA2.0 Gu et al., (2023). Their experiments demonstrated that these models are prone to generating biased content. Deng et al., (2022) focused on detecting offensive language in Chinese, proposing the Chinese Offensive Language Detection (COLD) benchmark, which includes a dataset and a baseline detector for identifying offensive language. While these studies primarily focused on pre-trained models and did not investigate stereotypes in commercial Chinese-based language models, they contributed to constructing social group categories within Chinese society. As we detail next, we combined the social groups identified by Huang and Xiong, (2024) and Zhao et al., (2023) to create a new taxonomy comprising 13 categories relevant to Chinese society (e.g., socioeconomic status, diseases).
@@
In a previous work Liu et al., (2024), we investigated auto-completion moderation policies in two major Western and Chinese search engines (Google and Baidu), examining stereotype moderation in commercial language technology applications in a comparative fashion. However, the social groups used in this study were adopted directly from U.S.-centric research and focused only on search engines, limiting their applicability to Chinese contexts. Here, we extend those analyses by (i) including Large Language Models in the study and (ii) investigating stereotypes and negative views about social groups pertinent to Chinese society.
@@ 3.1 Chinese social groups
To collect language models’ and search engines’ views on Chinese society, we combined the social groups described in Huang and Xiong, (2024) and Zhao et al., (2023). Huang and Xiong, (2024) identified different types of biases in Chinese culture: categories such as Disability, Disease, Ethnicity, and Gender were extracted from China’s “Employment Promotion Law” (Brown,, 2009), while additional categories like Age, Education, and Sexual Orientation were sourced from social media discussions on “Weibo” and “Zhihu” [^1] Weibo, often referred to as the “Chinese Twitter”, and “Zhihu”, similar to “Quora”, are popular platforms in China for sharing knowledge and engaging in discussions.. They also employed “CNKI” [^2] https://www.cnki.net/, a Chinese knowledge resource that includes journals, theses, conference papers, and books, to review qualitative and quantitative studies on these biases. This led to the creation of a dataset covering stereotypes and societal biases in 14 social dimensions and 100k related questions about Chinese culture and values, constructed with a semi-automatic approach employing annotators and generative language models, finally validated by human experts.
Zhao et al., (2023) followed the explicit bias specifications category from (Caliskan et al.,, 2017; Lauscher et al.,, 2020) to define four bias categories in Chinese: Age, Appearance, Gender, and Sexual Orientation.
After combining the aforementioned resources, one of the authors, who is a Chinese national and native speaker, selected unique social groups from both datasets based on the following principles:
@@
  • Remove social groups that were too time-specific like “COVID-19 patients”, “Post-2011s generation” and “Post-2012s generation”.
  • Merge similar groups. For instance, in the classification of socioeconomic status, common social groups like “Students from low-income families” and “Friends from low-income families” are consolidated under the representative term “People from low-income families”.
@@
The resulting dataset contains 240 social groups across 13 categories. We provide the breakdown in Table 1 along with some examples in English and Chinese. The full code and data to reproduce our results are publicly available in the repository associated with this paper [^3] https://github.com/leoleepsyche/stereotypes_in_search_engines-and-Chinese-LLMs.
@@ 3.2.1 Baidu
We collected search auto-completion data from Baidu, the largest online search engine in China (Zhang,, 2020). We will refer to it as a “model” in the rest of the text, to keep consistency with the LLMs that we describe next. According to the documentation of Baidu Baike, auto-completions from Baidu are based on hundreds of millions of user search terms history every data. [^4] https://baike.baidu.com/item/{CJK*}UTF8gbsn百度下拉/7139864
@@ 3.2.2 Ernie and Qwen
We collected data from the two main commercial Large Language Models in the Chinese market: Alibaba’s Tongyi Qianwen (Qwen) and Baidu’s Ernie. We chose these models as they are the most accessible and widely used modern LLMs in Chinese society, with both companies stating that they will be integrated into various projects in the future, making it worthwhile to explore potential stereotypes or biases in these models (Guo et al.,, 2023).
Qwen, released by Alibaba Cloud, consists of a series of models that are suitable for a wide range of NLP tasks in the Chinese language. For example, Qwen-72B was trained on an astounding 3 trillion tokens of data. The training procedure includes pre-training on large corpora such as Chinese, English, multilingual texts, code, and mathematics, covering general and professional fields. [^5] https://www.alibabacloud.com/en/solutions/generative-ai/qwen?_p_lc=1. Similarly, Ernie is a series of powerful models developed by Baidu based on the ERNIE (Enhanced Representation through Knowledge Integration) and PLATO (Pre-trained Dialogue Generation Model) series models. According to public information, the training corpus consists of a 4TB dataset spanning 11 categories, including search, web, novel, law, and so on, which is currently the largest Chinese pre-training corpus [^6] https://sh-tsang.medium.com/brief-review-ernie-3-0-large-scale-knowledge-enhanced-pre-training-for-language-understanding-7b0f777b19be.
Both series models offer APIs to interact with their functions. We chose the qwen-turbo of Qwen [^7] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/xlmokikxe and the ernie-turbo [^8] https://help.aliyun.com/zh/dashscope/developer-reference/api-details?spm=a2c4g.11186623.0.0.b22a3116jp5wvR version for Ernie. To obtain completion data from these models, we followed an approach similar to Baidu’s search engine based on previous work (Busker et al.,, 2023). We used the ChatCompletion functions in the models’ API setting the temperature parameter to the minimum, to ensure a deterministic output. Specifically, we set the temperature to 0.01, the max_output_token to 300 and the top_p to 0.99. For Ernie, the max_output_token can range from 2 to 2048, while for Qwen Turbo, it is up to 1500. For the temperature and top_p parameters, Ernie allows values in the range $(0,1.0]$ and $[0,1.0]$, while Qwen’s valid ranges are $[0,2)$ and $(0,1)$.
@@
  • system: {CJK*}UTF8gbsn”你是一个句子补全助手。请补全下列句子中的缺失部分,并生成10个可能的补全选项,请将结果以JSON格式输出.” (English: ”You are a sentence completion assistant. Please complete the missing parts of the following sentence and generate 10 possible completion options. Please output the results in JSON format”).
The user prompts correspond to the templates described in Table 2. We required the models to return 10 completions as this is the maximum number of results returned by Baidu. We performed 240 (groups) · 6 (templates) = 1440 queries for both Ernie and Qwen. We notice that the models did not necessarily return 10 results are requested in the instructions. The total number of completions is 14302 for Qwen and 14131 for Ernie.

### v3

@@ Abstract
Large Language Models (LLMs) and search engines have the potential to perpetuate biases and stereotypes by amplifying existing prejudices in their training data and algorithmic processes, thereby influencing public perception and decision-making. While most work has focused on Western-centric AI technologies, we examine social biases embedded in prominent Chinese-based commercial tools, the main search engine Baidu and two leading LLMs, Ernie and Qwen. Leveraging a dataset of 240 social groups across 13 categories describing Chinese society, we collect over 30k views encoded in the aforementioned tools by prompting them to generate candidate words describing these groups. We find that language models exhibit a broader range of embedded views compared to the search engine, although Baidu and Qwen generate negative content more often than Ernie. We also observe a moderate prevalence of stereotypes embedded in the language models, many of which potentially promote offensive or derogatory views. Our work highlights the importance of prioritizing fairness and inclusivity in AI technologies from a global perspective.
@@ 1 Introduction
The advent of novel Large Language Models (LLMs) has revolutionized the field of natural language processing (NLP), offering unprecedented potential for a wide range of applications, from machine translation and sentiment analysis to conversational agents and content generation Zhu et al., 2024b ; Zhang et al., (2024); Deng et al., (2023). LLMs have demonstrated an impressive ability to understand and generate human-like text, making them invaluable tools in both academic research and commercial settings Ziems et al., (2024); Minaee et al., (2024). However, along with their potential, these models also carry significant risks. As language models (LMs) become increasingly integrated into various applications, concerns about their reliability, ethical use, and potential for misuse have grown, prompting a growing body of studies first related to early generation language models May et al., (2019); Nadeem et al., (2021) and more recently to LLMs Choenni et al., (2021); Navigli et al., (2023); Busker et al., (2023); Deshpande et al., (2023); Gallegos et al., (2024). In addition, the opaque nature of these models often makes it difficult to understand their inference processes, raising questions about accountability and transparency Bender et al., (2021); Liao and Wortman Vaughan, (2024).
@@
One pressing issue associated with this technology is the presence of biases and stereotypes embedded within language models, which can stem from the data they are trained on, thus reflecting the prejudices and inequalities prevalent in broader cultural and social contexts Liang et al., (2020); Kotek et al., (2023). These biases can manifest in harmful ways, such as reinforcing stereotypes or supporting discriminatory decisions, thus perpetuating social injustices Weidinger et al., (2021); Nogara et al., (2025); Bär et al., (2024). Researchers have developed various methods to measure and mitigate these biases, proposing datasets and metrics to evaluate the models’ fairness Zmigrod et al., (2019); Webster et al., (2020); Liang et al., (2020). Despite these efforts, significant challenges remain, particularly when addressing proprietary models where access to internal mechanisms remains restricted Huang et al., (2024). This limitation hampers the effectiveness of many bias detection and mitigation techniques, underscoring the necessity for further innovation and scrutiny in this area.
@@
Most research on bias and fairness in LLMs has focused on English and other widely spoken languages, leaving a gap in our understanding of how these models perform in different linguistic and cultural contexts Ducel et al., (2023); Yang et al., (2024). Chinese, natively spoken by over 1.35 billion individuals (${\sim}17\%$ of the global population), has received comparatively little attention from the scientific community Xu et al., (2024). This oversight is critical, given the unique cultural, social, and linguistic characteristics of the Chinese language. Most of the current LLMs are primarily trained on English corpora, which inherently reflect Western cultural values Wang et al., 2024a ; Naous et al., (2024); Li et al., (2024); Zhu et al., 2024a . However, when such models are employed in non-Western settings, biases and distortions may arise since language and culture are intrinsically linked with each other, as language reflects and conveys cultural values, norms, and perspectives woven into society. For example, LLMs may generate responses that assume culture-specific norms, such as referencing holidays like Thanksgiving, which are not celebrated globally. In contrast, Ernie, developed by Baidu, demonstrates a stronger alignment with Chinese cultural contexts, providing culturally appropriate responses when prompted in Chinese, such as references to Lunar New Year and the Qixi Festival (Chinese Valentine’s Day) Wang et al., 2024a . Likewise, ChatGPT has demonstrated limitations in accurately interpreting medical knowledge specific to the Chinese population, particularly concerning Traditional Chinese Medicine (TCM). It has been shown that Chinese LLMs such as Qwen and Ernie outperform Western models on TCM-related questions, achieving an average accuracy of 78.4% compared to 35.9% for Western models Zhu et al., 2024a . Addressing cultural specificities is therefore essential for developing fair and unbiased LLMs that can serve diverse populations effectively and ethically Zhao et al., (2023); Huang and Xiong, (2024). Consequently, our study aims to investigate the biases present in two leading Chinese LLMs, namely Ernie and Qwen, as well as in Baidu, the most popular search engine used in China; we include the latter following previous work Choenni et al., (2021); Busker et al., (2023); Liu et al., (2024) that highlighted the role of online search engines in potentially perpetuating cultural biases and stereotypes. Leveraging a dataset of social groups and categories pertaining to Chinese society, we use an empirical approach to study the social views encoded in the models. Following Busker et al., (2023), we probe the models with multiple questions, to elicit the views that are encoded about different social groups. We then study the completions obtained in terms of diversity, negativity, and alignment. We use this approach as the internal models’ probability for predicted tokens is not available, and existing metrics for measuring biases associated with certain words cannot be applied Kurita et al., (2019); Gallegos et al., (2024). Our work contributes to the literature on fairness and inclusivity in AI-generated content by systematically evaluating Chinese-based LLMs (Ernie and Qwen) and comparing them with a Chinese search engine (Baidu). To the best of our knowledge, this is the first comparative study that investigates biases across multiple proprietary Chinese language AI technologies. Our analysis reveals significant diversity gaps and varying levels of bias, highlighting the risks associated with biased outputs concerning different social groups. Additionally, our study contributes to understanding and mitigating biases present in Chinese-based LLMs. Addressing these biases is not only a technical challenge but also an ethical responsibility, as it directly affects how different social groups are perceived and represented in society.
@@
The manuscript is organized as follows: in the next section, we review related work; we then describe the data collection and methodologies used in our analyses; next, we present the results of our analyses; finally, we discuss our findings and their implications, mention limitations, and draw conclusions.
@@ 2 Related Work
In recent years, various studies have introduced datasets, methods and metrics to quantify social bias and stereotypes in pre-trained early-generation LMs and LLMs Caliskan et al., (2017); Zhao et al., (2018); May et al., (2019); Kurita et al., (2019); Liang et al., (2020); Barikeri et al., (2021); Wald and Pfahler, (2023). However, these efforts face challenges with most proprietary models, as access to internal model information (e.g., probability of predicted tokens and internal embeddings) is restricted Huang et al., (2023), thereby limiting the application of metrics such as the Log Probability Bias Score (LPBS) Kurita et al., (2019) and the Sentence Embedding Association Test (SEAT) May et al., (2019). In addition, the literature advises to exercise caution with embedding- and probability-based metrics  Gallegos et al., (2024). In the following, we focus on contributions more directly related to our work and refer readers to  Gallegos et al., (2024) for a more comprehensive review of the literature on this topic.
@@
Choenni et al., (2021) introduced a method for eliciting stereotypes by retrieving salient attributes associated with social groups through the auto-completion functionality in search engines such as Google, Yahoo, and DuckDuckGo. They then analyzed the predictions of pre-trained language models to assess how many of these stereotypical attributes were encoded within the models. Their study also highlighted the evolution of stereotypes with modifications in training data during model fine-tuning stages. Building on this work, Busker et al., (2023) conducted an empirical study to explore stereotypical behavior in ChatGPT, focusing on U.S.-centric social groups. They presented stereotypical prompts in six formats (e.g., questions and statements) and across nine different social group categories (e.g., age, country, and profession). For each prompt, they asked ChatGPT to fill in a masked word and mapped the completions to sentiment levels to measure stereotypical behavior.
@@
To date, only a few studies have addressed biases and stereotypes in Chinese-based language models. Huang and Xiong, (2024) extended the Biases Benchmark (BBQ) dataset Parrish et al., (2022) from U.S. English-speaking contexts to Chinese society, creating the Chinese Biases Benchmark (CBBQ) dataset with contributions from human experts and generative language models such as GPT-4. This dataset encompasses a wide range of biases and stereotypes related to Chinese culture and content. Similarly, Zhao et al., (2023) introduced the CHBias dataset, which includes biases related to ageism, appearance, and other factors, for evaluating and mitigating biases in Chinese conversational models like CDial-GPT Wang et al., (2020) and EVA2.0 Gu et al., (2023). Their experiments demonstrated that these models are prone to generating biased content. Deng et al., (2022) focused on detecting offensive language in Chinese, proposing the Chinese Offensive Language Detection (COLD) benchmark, which includes a dataset and a baseline detector for identifying offensive language. These studies primarily focused on pre-trained models and did not investigate stereotypes in commercial Chinese-based language models, they contributed to constructing social group categories within Chinese society. As we detail next, we combined the social groups identified by Huang and Xiong, (2024) and Zhao et al., (2023) to create a new taxonomy comprising 13 categories relevant to Chinese society (e.g., socioeconomic status, diseases).
@@
In a previous work Liu et al., (2024) we investigated auto-completion moderation policies in two major Western and Chinese search engines (Google and Baidu), examining stereotype moderation in commercial language technology applications in a comparative fashion. However, the social groups used in that study were drawn directly from U.S.-centric research and focused only on search engines, limiting their applicability to Chinese contexts. Here, we extend those analyses by (i) including Large Language Models in the study and (ii) investigating stereotypes and negative views about social groups pertinent to Chinese society.
@@ 3 Materials and Methods
Portions of this text were previously published as part of a preprint (https://arxiv.org/pdf/2408.15696).
@@ 3.1 Chinese social groups
To collect language models’ and search engines’ views on Chinese society, we combined the social groups described in Huang and Xiong, (2024) and Zhao et al., (2023). Huang and Xiong, (2024) identified different types of biases in Chinese culture: categories such as Disability, Disease, Ethnicity, and Gender were extracted from China’s “Employment Promotion Law” Brown, (2009) while additional categories like Age, Education, and Sexual Orientation were sourced from social media discussions on “Weibo” and “Zhihu” [^1] Weibo, often referred to as the “Chinese Twitter”, and “Zhihu”, similar to “Quora”, are popular platforms in China for sharing knowledge and engaging in discussions.. They also employed “CNKI” [^2] https://www.cnki.net/, a Chinese knowledge resource that includes journals, theses, conference papers, and books, to review qualitative and quantitative studies on these biases. This led to the creation of a dataset covering stereotypes and societal biases in 14 social dimensions and 100k related questions about Chinese culture and values, constructed with a semi-automatic approach employing annotators and generative language models, finally validated by human experts.
Zhao et al., (2023) followed the explicit bias specifications category from Caliskan et al., (2017); Lauscher et al., (2020) to define four bias categories in Chinese: Age, Appearance, Gender, and Sexual Orientation.
After merging the social groups from the specified sources, one of the authors, who is a Chinese national and native speaker, selected specific social groups based on the following principles:
@@
  • Remove social groups that were highly time-specific, such as “COVID-19 patients”, “Post-2011s generation” and “Post-2012s generation”.
  • Merge similar groups where appropriate. For instance, in the classification of socioeconomic status, common social groups like “Students from low-income families” and “Friends from low-income families” were consolidated under the representative term “People from low-income families”.
@@
The resulting dataset contains 240 social groups across 13 categories. We provide the breakdown in Table 1 along with some examples in English and Chinese. The full code and data to reproduce our results are publicly available in the repository associated with this paper [^3] https://doi.org/10.5281/zenodo.14148258. The workflow of our analyses is shown in Figure 1.
@@ 3.2.1 Baidu
We collected search auto-completion data from Baidu, the largest online search engine in China Zhang, (2020). We will refer to it as a “model” in the rest of the text, to keep consistency with the LLMs that we describe next. According to the documentation of Baidu Baike, auto-completions from Baidu are based on hundreds of millions of user search terms per day. [^4] https://baike.baidu.com/item/{CJK*}UTF8gbsn百度下拉/7139864
@@
We also considered including English templates for Baidu but found that the majority of the queries would not return auto-completion results. This limitation is most likely due to the fact that Baidu’s user base is Chinese, and most queries are made in Chinese language. Therefore, we only considered templates in the Chinese language.
@@ 3.2.2 Ernie and Qwen
We collected data from the two main commercial LLMs in the Chinese market: Alibaba’s Tongyi Qianwen (Qwen) and Baidu’s Ernie. We chose these models as Alibaba and Baidu are the most popular AI companies in Chinese society, making it crucial to examine potential stereotypes or biases in these models Guo et al., (2023).
Qwen, released by Alibaba Cloud, consists of a series of models that are suitable for a wide range of NLP tasks in the Chinese language, including Qwen-72B which was trained on 3 trillion tokens of data comprising Chinese, English and multilingual text as well as code, mathematics, and other fields. [^5] https://www.alibabacloud.com/en/solutions/generative-ai/qwen?_p_lc=1. Similarly, Ernie is a series of powerful models developed by Baidu based on the ERNIE (Enhanced Representation through Knowledge Integration) and PLATO (Pre-trained Dialogue Generation Model) models. According to public information, the training corpus consists of a 4TB dataset spanning 11 categories, including search, web, novel, law, and so on, which is currently the largest Chinese pre-training corpus [^6] https://sh-tsang.medium.com/brief-review-ernie-3-0-large-scale-knowledge-enhanced-pre-training-for-language-understanding-7b0f777b19be.
Both models offer APIs to interact with their functions. We chose the qwen-turbo for Qwen [^7] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/xlmokikxe and the ernie-turbo [^8] https://help.aliyun.com/zh/dashscope/developer-reference/api-details?spm=a2c4g.11186623.0.0.b22a3116jp5wvR version for Ernie. We used the ChatCompletion functions in the models’ API setting the temperature parameter to the minimum to ensure a deterministic output. Specifically, we set the temperature to 0.01, the max_output_token to 300 and the top_p to 0.99. For Ernie, the max_output_token can range from 2 to 2048, while for Qwen Turbo it is up to 1500. For the temperature and top_p parameters, Ernie allows values in the range $(0,1.0]$ and $[0,1.0]$, while Qwen’s valid ranges are $[0,2)$ and $(0,1)$.
@@
  • system: {CJK*}UTF8gbsn“你是一个句子补全助手。请补全下列句子中的缺失部分,并生成10个可能的补全选项,请将结果以JSON格式输出.” (English: “You are a sentence completion assistant. Please complete the missing parts of the following sentence and generate 10 possible completion options. Please output the results in JSON format”).
[…]